OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find f(-1) + 5g(-3).

OpenStudy (anonymous):

go to: http://www.wolframalpha.com/, and enter your numbers. it will give you answer with steps and graphs!

OpenStudy (anonymous):

thats not helping i like it worked out the way @cody_123 does it

OpenStudy (anonymous):

\[(-1)^2+1+5(3*-3+1)\]

OpenStudy (anonymous):

-38

OpenStudy (anonymous):

thanks @cody_123 I love how u do math

OpenStudy (anonymous):

ali_hassan Best Response Medals 1 f(-1)=(-1)^2+1=2 g(-3)=3(-3)+1=-9+1=-8 5g(-3)=5(-8)=-40 f(-1)+5g(-3)=2-40=-38

OpenStudy (anonymous):

i found thi for u

OpenStudy (anonymous):

i did a google search with your problem

OpenStudy (anonymous):

some of my problems dont go in thats why i chose this place